OpenStudy (anonymous):

how do you solve x+y=4 x-y=-8 solving it as a linear equation using addition?

OpenStudy (anonymous):

x and y are 2 for x-y=4 x is 10 and y is 2 for x-y=8

OpenStudy (anonymous):

simultaneous equations add both equations to get rid of the y's in this question 2x = -4 x=-2

OpenStudy (anonymous):

Subs in any of the two equations to get y -2+y=4 in first equation y=6